HIGH-RESOLUTION SPECTRA OF MOLECULAR ANIONS: ANOMALOUS ASYMMETRY SPLITTING IN HNO^{−}

\underline{JOHN W. FARLEY}, Department of Physics, University of Nevada, Las Vegas, NV 89154; and DANIEL C. COWLES, Air Liquide Laboratories, 5-9-9, Tokodai, Tsukuba-shi, Ibaraki-Pref. 300-26, Japan.

Extensive measurements have been made of the infrared vibrational-rotational spectra of the H^{14}NO^{−} and H^{15}NO^{−} anions. The spectra were obtained using autodetachment spectroscopy, in which vibrational-rotational transitions are excited using a coaxial ion beam/infrared laser beam spectrometer. The vibrationally excited ions autodetach, producing a fast neutral which is detected. A total of 413 transitions have been measured in the fundamental N-H stretch. The measurements have sub-Doppler resolution, enabling the measurement of the asymmetry splitting. The asymmetry splitting, and rotational constants returned by a least-squares fit to the spectra, manifest an anomalously large dependence of the rotational constants upon isotopic substitution. The spectrum may be perturbed by a Coriolis interaction. Comparison is made with a theoretical calculation ^{a} of the geometry of this species.

aK. A. Robins, J. W. Farley, and J. L. Toto, J. Chem. Phys. 99, 9770 (1993).